# DCPI-Depth: Explicitly Infusing Dense Correspondence Prior to Unsupervised Monocular Depth Estimation

### v1

@@ Abstract
There has been a recent surge of interest in learning to perceive depth from monocular videos in an unsupervised fashion. A key challenge in this field is achieving robust and accurate depth estimation in challenging scenarios, particularly in regions with weak textures or where dynamic objects are present. This study makes three major contributions by delving deeply into dense correspondence priors to provide existing frameworks with explicit geometric constraints. The first novelty is a contextual-geometric depth consistency loss, which employs depth maps triangulated from dense correspondences based on estimated ego-motion to guide the learning of depth perception from contextual information, since explicitly triangulated depth maps capture accurate relative distances among pixels. The second novelty arises from the observation that there exists an explicit, deducible relationship between optical flow divergence and depth gradient. A differential property correlation loss is, therefore, designed to refine depth estimation with a specific emphasis on local variations. The third novelty is a bidirectional stream co-adjustment strategy that enhances the interaction between rigid and optical flows, encouraging the former towards more accurate correspondence and making the latter more adaptable across various scenarios under the static scene hypotheses. DCPI-Depth, a framework that incorporates all these innovative components and couples two bidirectional and collaborative streams, achieves state-of-the-art performance and generalizability across multiple public datasets, outperforming all existing prior arts. Specifically, it demonstrates accurate depth estimation in texture-less and dynamic regions, and shows more reasonable smoothness.
@@ I Introduction
Monocular depth estimation, a crucial research field in computer vision and robotics, has applications across various domains, such as autonomous driving [1], augmented reality [2], and embodied artificial intelligence [3]. It provides agents with powerful environmental perception capabilities, enabling robust ego-motion estimation and 3D geometry reconstruction [4]. Early monocular depth estimation approaches [5, 6, 7], developed based on supervised learning, typically require a large amount of well-annotated, per-pixel depth ground truth (generally acquired using high-precision LiDARs) for model training [8, 9]. Nevertheless, collecting and labeling such data is tedious and costly [10, 8]. Thus, the practical use of these supervised approaches remains limited.
In recent years, un/self-supervised monocular depth estimation has garnered significant attention [4, 10, 11, 12, 13, 14]. Such approaches obviate the need for extensive depth ground truth by leveraging either stereo image pairs [15, 16] or monocular videos to jointly learn depth and ego-motion estimation [4]. Specifically, given the target and source images, the estimated depth map (at the target view) and camera ego-motion are employed to warp the source image into the target view. The depth estimation network (hereafter referred to as DepthNet) and the pose estimation network (hereafter referred to as PoseNet) are then jointly trained in an un/self-supervised manner by minimizing the photometric loss, which measures the consistency between the reconstructed and original target images [4, 10, 13].
@@ II-A Supervised Monocular Depth Estimation
Supervised monocular depth estimation approaches require depth ground truth for model training. As the first attempt, the study [5] proposes a coarse-to-fine architecture and a scale-invariant loss function to perceive depth from a single image. Subsequent research focuses mainly on improving depth estimation performance by exploring more intricate network architectures [19, 6, 7] or developing new loss functions [20, 21, 22]. For example, in studies [19] and [22], monocular depth estimation was reformulated as a per-pixel classification task, where depth ranges instead of exact depth values are predicted. Furthermore, to combine the benefits of both regression and classification tasks, the study [23] redefines this problem as a per-pixel classification-regression task. In [24], multi-scale guidance layers are introduced to establish connections between intermediate-layer features and the final depth map. In [25], a vision Transformer (ViT)-based architecture was developed to capture long-range correlations in depth estimation. More recently, Depth Anything [26] has demonstrated impressive performance, primarily due to its powerful backbone (a ViT-based vision foundation model) that is capable of extracting general-purpose, informative deep features. It first reproduces a MiDaS-based [27] teacher model with pre-trained weights from DINOv2 [28], and then utilizes the teacher’s predictions as pseudo-labels to train a student model on large-scale unlabeled data. While these methods have shown promising results, the need for extensive ground-truth depth data for supervised model training imposes considerable costs, thus limiting their wider applicability, especially in autonomous driving. Therefore, un/self-supervised monocular depth estimation approaches require greater attention.
@@ II-B Un/Self-Supervised Monocular Depth Estimation
To liberate monocular depth estimation from dependence on extensive ground-truth data, un/self-supervised approaches [15, 4, 10, 13, 14] have emerged as visible alternatives. These methods typically utilize estimated depth to establish a differentiable warping between two images and employ photometric loss as the supervisory signal [29, 4]. The study [29] represents the first reported attempt to learn monocular depth estimation from stereo image pairs within a self-supervised framework. Subsequently, the study [4] extends this approach by coupling the learning of depth and ego-motion estimation from monocular videos. However, un/self-supervised methods often face challenges with independently moving objects and preserving clear object boundaries, due to multi-view ambiguities [13]. Therefore, in [10], a minimum reprojection loss and an auto-masking technique were introduced to exclude such regions during model training, significantly improving depth estimation performance. Building upon these prior arts, several studies have explored more sophisticated network architectures [30, 31, 32, 33, 34, 14, 35] for improved performance. Others have incorporated multi-task learning, such as optical flow estimation [36, 37, 38] and semantic segmentation [39, 40], to address the challenges with dynamic objects.
@@ III-C Differential Property Correlation Loss
${}^{\boldsymbol{{\rm t}}}\boldsymbol{F}^{O}_{t\to s}=\boldsymbol{F}^{O}_{t\to s}-{{}^{\boldsymbol{{\rm R}}}\boldsymbol{F}^{R}_{t\to s}},$ (9)
where ${}^{\boldsymbol{{\rm t}}}\boldsymbol{F}^{O}_{t\to s}$ denotes the translational optical flow, and ${}^{\boldsymbol{{\rm R}}}\boldsymbol{F}^{R}_{t\to s}$, the rotational rigid flow, is generated using the estimated $\boldsymbol{R}$ and a translation vector of zeros $\boldsymbol{0}$ as follows:
$\begin{bmatrix}{}^{\boldsymbol{{\rm R}}}\boldsymbol{F}^{R}_{t\to s}(\boldsymbol{p}_{t})\\ 0\end{bmatrix}+\tilde{\boldsymbol{p}}_{t}\sim\boldsymbol{K}\begin{bmatrix}\boldsymbol{R}&\boldsymbol{0}\end{bmatrix}\begin{bmatrix}\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})\boldsymbol{K}^{-1}\tilde{\boldsymbol{p}}_{t}\\ 1\end{bmatrix}.$ (10)
@@
$\begin{bmatrix}{}^{\boldsymbol{{\rm t}}}\boldsymbol{F}_{t\to s}^{O}(\boldsymbol{p}_{t})\\ 0\end{bmatrix}$ $=\displaystyle\frac{\boldsymbol{K}\begin{bmatrix}\boldsymbol{I}&\boldsymbol{t}\end{bmatrix}\begin{bmatrix}\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})\boldsymbol{K}^{-1}\tilde{\boldsymbol{p}}_{t}\\ 1\end{bmatrix}}{\boldsymbol{D}^{C}_{s}(\boldsymbol{p}_{s})}-\tilde{\boldsymbol{p}}_{t}$ (11)
@@
$\nabla\cdot{}^{\boldsymbol{{\rm t}}}\boldsymbol{F}_{t\to s}^{O}(\boldsymbol{p}_{t})=\nabla\cdot{}\left(\frac{t_{3}}{\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})-t_{3}}\boldsymbol{q}_{t}\right)$
$=$ $\boldsymbol{q}_{t}\cdot{}\nabla\frac{t_{3}}{\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})-t_{3}}+\frac{t_{3}}{\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})-t_{3}}\nabla\cdot{}\boldsymbol{p}_{t},$ (12)
@@
$\underbrace{\frac{\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})-t_{3}}{t_{3}}\nabla\cdot{}^{\boldsymbol{{\rm t}}}\boldsymbol{F}_{t\to s}^{O}(\boldsymbol{p}_{t})-\nabla\cdot{}\boldsymbol{p}_{t}}_{\boldsymbol{C}^{F}(\boldsymbol{p}_{t})}$
$=$ $\underbrace{-\frac{\boldsymbol{q}_{t}}{\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})-t_{3}}\cdot{}\nabla\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})}_{\boldsymbol{C}^{D}(\boldsymbol{p}_{t})}.$ (14)
@@
TABLE I: Quantitative comparison with SoTA networks on the KITTI [45] dataset. The best results are shown in bold symbol. The symbols $\uparrow$ and $\downarrow$ indicate that higher and lower values correspond to better performance, respectively. “M” denotes training with monocular video sequences.
Monodepth2[10] 2019 192 $\times$ 640 M 0.115 0.903 4.863 0.193 0.877 0.959 0.981
HR-Depth[43] 2020 192 $\times$ 640 M 0.109 0.792 4.632 0.185 0.884 0.962 0.983
DIFFNet[8] 2021 192 $\times$ 640 M 0.102 0.764 4.483 0.180 0.890 0.964 0.983
MonoViT-tiny[9] 2022 192 $\times$ 640 M 0.102 0.733 4.459 0.177 0.895 0.965 0.984
Swin-Depth[46] 2023 192 $\times$ 640 M 0.106 0.739 4.510 0.182 0.890 0.964 0.984
Lite-Mono[14] 2023 192 $\times$ 640 M 0.107 0.765 4.561 0.183 0.886 0.963 0.983
Lite-Mono-8M[14] 2023 192 $\times$ 640 M 0.101 0.729 4.454 0.178 0.897 0.965 0.983
Dynamo-Depth[47] 2023 192 $\times$ 640 M 0.112 0.758 4.505 0.183 0.873 0.959 0.984
DaCCN[35] 2023 192 $\times$ 640 M 0.099 0.661 4.316 0.173 0.897 0.967 0.985
SENSE[48] 2023 192 $\times$ 640 M 0.104 0.693 4.294 0.177 0.894 0.965 0.984
AQUANet[49] 2024 192 $\times$ 640 M 0.105 0.621 4.227 0.179 0.889 0.964 0.984
DCPI-Depth (Ours) - 192 $\times$ 640 M 0.095 0.662 4.274 0.170 0.902 0.967 0.985
HR-Depth[43] 2020 320 $\times$ 1024 M 0.106 0.755 4.472 0.181 0.892 0.966 0.984
@@ III-D Bidirectional Stream Co-Adjustment Strategy
In static regions, where optical flow and rigid flow should ideally be identical, (16) allows both flows to co-adjust without compromising photometric consistency and encourages rigid flow towards more accurate correspondences. On the other hand, in dynamic regions where optical flow and rigid flow differ significantly, (16) encourages the unfrozen FlowNet to generate “quasi” rigid flow, eliminating independent apparent motions that can mislead DepthNet, thereby retaining only the apparent motions consistent with the static scene hypothesis. This enables our CGDC and DPC losses to be effectively applied across the entire image, irrespective of static and dynamic regions. The generated “quasi” rigid flow produced by the adjusted FlowNet can, in turn, optimize depth estimation in dynamic regions via our developed CGDC and DPC losses. We conduct an ablation study in Sect. IV-E to demonstrate the effectiveness of this strategy.
@@
TABLE II: Quantitative comparison with SoTA networks on the KITTI [45] dataset using the improved KITTI ground truth provided in [50]. The best results are shown in bold symbol. The symbols $\uparrow$ and $\downarrow$ indicate that higher and lower values correspond to better performance, respectively. “M” denotes training with monocular video sequences, and “MS” denotes training with both monocular video sequences and stereo image pairs. indicates the results achieved using the same weights in Table I for consistency.

### v2

@@ Abstract
There has been a recent surge of interest in learning to perceive depth from monocular videos in an unsupervised fashion. A key challenge in this field is achieving robust and accurate depth estimation in challenging scenarios, particularly in regions with weak textures or where dynamic objects are present. This study makes three major contributions by delving deeply into dense correspondence priors to provide existing frameworks with explicit geometric constraints. The first novelty is a contextual-geometric depth consistency loss, which employs depth maps triangulated from dense correspondences based on estimated ego-motion to guide the learning of depth perception from contextual information, since explicitly triangulated depth maps capture accurate relative distances among pixels. The second novelty arises from the observation that there exists an explicit, deducible relationship between optical flow divergence and depth gradient. A differential property correlation loss is, therefore, designed to refine depth estimation with a specific emphasis on local variations. The third novelty is a bidirectional stream co-adjustment strategy that enhances the interaction between rigid and optical flows, encouraging the former towards more accurate correspondence and making the latter more adaptable across various scenarios under the static scene hypotheses. DCPI-Depth, a framework that incorporates all these innovative components and couples two bidirectional and collaborative streams, achieves state-of-the-art performance and generalizability across multiple public datasets, outperforming all existing prior arts. Specifically, it demonstrates accurate depth estimation in texture-less and dynamic regions, and shows more reasonable smoothness. Our source code will be publicly available at mias.group/DCPI-Depth upon publication.
@@ I Introduction
Monocular depth estimation, a crucial research field in computer vision and robotics, has applications across various domains, such as autonomous driving [1], augmented reality [2], and embodied artificial intelligence [3]. It provides agents with powerful environmental perception capabilities, enabling robust ego-motion estimation and 3D geometry reconstruction [4]. Early monocular depth estimation approaches [5, 6], developed based on supervised learning, typically require a large amount of well-annotated, per-pixel depth ground truth (generally acquired using high-precision LiDARs) for model training [7, 8]. Nevertheless, collecting and labeling such data is tedious and costly [9, 7]. Thus, the practical use of these supervised approaches remains limited.
In recent years, un/self-supervised monocular depth estimation has garnered significant attention [4, 9, 10, 11, 12]. Such approaches obviate the need for extensive depth ground truth by leveraging either stereo image pairs [13, 14] or monocular videos to jointly learn depth and ego-motion estimation [4]. Specifically, given the target and source images, the estimated depth map (at the target view) and camera ego-motion are employed to warp the source image into the target view. The depth estimation network (hereafter referred to as DepthNet) and the pose estimation network (hereafter referred to as PoseNet) are then jointly trained in an un/self-supervised manner by minimizing the photometric loss, which measures the consistency between the reconstructed and original target images [4, 9, 11].
@@ II-A Supervised Monocular Depth Estimation
Supervised monocular depth estimation approaches require depth ground truth for model training. As the first attempt, the study [5] proposes a coarse-to-fine architecture and a scale-invariant loss function to perceive depth from a single image. Subsequent research focuses mainly on improving depth estimation performance by exploring more intricate network architectures [17, 6] or developing new loss functions [18, 19]. For example, in studies [17] and [19], monocular depth estimation was reformulated as a per-pixel classification task, where depth ranges instead of exact depth values are predicted. Furthermore, to combine the benefits of both regression and classification tasks, the study [20] redefines this problem as a per-pixel classification-regression task. In [21], multi-scale guidance layers are introduced to establish connections between intermediate-layer features and the final depth map. In [22], a vision Transformer (ViT)-based architecture was developed to capture long-range correlations in depth estimation. More recently, Depth Anything [23] has demonstrated impressive performance, primarily due to its powerful backbone (a ViT-based vision foundation model) that is capable of extracting general-purpose, informative deep features. It first reproduces a MiDaS-based [24] teacher model with pre-trained weights from DINOv2 [25], and then utilizes the teacher’s predictions as pseudo-labels to train a student model on large-scale unlabeled data. Building on single-image depth estimation, monocular video depth estimation incorporates both temporal and geometric consistencies [2]. Representative works, such as [2] and [26], pioneer the use of correspondences and camera poses to enforce inter-frame depth consistency in 3D space, which inspires us to further explore the way of infusing dense correspondence prior to guide the unsupervised learning for monocular depth estimation.
@@ II-B Un/Self-Supervised Monocular Depth Estimation
To liberate monocular depth estimation from dependence on extensive ground-truth data, un/self-supervised approaches [13, 4, 9, 11, 12] have emerged as visible alternatives. These methods typically utilize estimated depth to establish a differentiable warping between two images and employ photometric loss as the supervisory signal [27, 4]. The study [27] represents the first reported attempt to learn monocular depth estimation from stereo image pairs within a self-supervised framework. Subsequently, the study [4] extends this approach by coupling the learning of depth and ego-motion estimation from monocular videos. However, un/self-supervised methods often face challenges with independently moving objects and preserving clear object boundaries, due to multi-view ambiguities [11]. Therefore, in [9], a minimum reprojection loss and an auto-masking technique were introduced to exclude such regions during model training, significantly improving depth estimation performance. Building upon these prior arts, several studies have explored more sophisticated network architectures [28, 29, 30, 12, 31] for improved performance. Others incorporate additional relevant tasks such as optical flow estimation [32, 33, 34] and semantic segmentation [35, 36] to enhance cross-task consistency or address the dynamic object challenge.
@@ III-C Differential Property Correlation Loss
$\boldsymbol{F}^{\rm Tra}_{t\to s}=\boldsymbol{F}^{O}_{t\to s}-{\boldsymbol{F}^{\rm{Rot}}_{t\to s}},$ (9)
where $\boldsymbol{F}^{\rm Tra}_{t\to s}$ denotes the translational optical flow, and $\boldsymbol{F}^{\rm Rot}_{t\to s}$, the rotational rigid flow, is generated using the estimated $\boldsymbol{R}$ and a translation vector of zeros $\boldsymbol{0}$ as follows:
$\begin{bmatrix}\boldsymbol{F}^{\rm Rot}_{t\to s}(\boldsymbol{p}_{t})\\ 0\end{bmatrix}+\tilde{\boldsymbol{p}}_{t}\sim\boldsymbol{K}\begin{bmatrix}\boldsymbol{R}&\boldsymbol{0}\end{bmatrix}\begin{bmatrix}\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})\boldsymbol{K}^{-1}\tilde{\boldsymbol{p}}_{t}\\ 1\end{bmatrix}.$ (10)
@@
$\begin{bmatrix}\boldsymbol{F}^{\rm Tra}_{t\to s}(\boldsymbol{p}_{t})\\ 0\end{bmatrix}$ $=\frac{\boldsymbol{K}\begin{bmatrix}\boldsymbol{I}&\boldsymbol{t}\end{bmatrix}\begin{bmatrix}\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})\boldsymbol{K}^{-1}\tilde{\boldsymbol{p}}_{t}\\ 1\end{bmatrix}}{\boldsymbol{D}^{C}_{s}(\boldsymbol{p}_{s})}-\tilde{\boldsymbol{p}}_{t}$ (11)
@@
$\nabla\cdot{}\boldsymbol{F}^{\rm Tra}_{t\to s}(\boldsymbol{p}_{t})=\nabla\cdot{}\left(\frac{t_{3}}{\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})-t_{3}}\boldsymbol{q}_{t}\right)$ (12)
$=$ $\boldsymbol{q}_{t}\cdot{}\nabla\frac{t_{3}}{\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})-t_{3}}+\frac{t_{3}}{\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})-t_{3}}\nabla\cdot{}\boldsymbol{p}_{t},$
@@
$\underbrace{\frac{\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})-t_{3}}{t_{3}}\nabla\cdot{}\boldsymbol{F}^{\rm Tra}_{t\to s}(\boldsymbol{p}_{t})-\nabla\cdot{}\boldsymbol{p}_{t}}_{\boldsymbol{C}^{F}(\boldsymbol{p}_{t})}$ (14)
$=$ $\underbrace{-\frac{\boldsymbol{q}_{t}}{\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})-t_{3}}\cdot{}\nabla\boldsymbol{D}^{C}_{t}(\boldsymbol{p}_{t})}_{\boldsymbol{C}^{D}(\boldsymbol{p}_{t})}.$
@@ III-D Bidirectional Stream Co-Adjustment Strategy
In static regions, optical flow and rigid flow should ideally be identical, as demonstrated in prior studies [32, 42]. Similar to [32], (16) allows both flows to adjust together without compromising photometric consistency. It particularly encourages rigid flow in the PCG stream towards more accurate correspondences. On the other hand, in dynamic regions, optical flow and rigid flow should differ significantly. However, the study [32] introduces an untrained FlowNet to form a joint learning framework, where both FlowNet and DepthNet are trained by minimizing the photometric loss. Directly minimizing (16) in this training paradigm can cause the DepthNet to be further misled in dynamic regions. In contrast, we decouple the FlowNet training from this joint learning framework, generating prior optical flow using a pre-trained FlowNet, which is updated solely by (16). As shown in Fig. 3, by ignoring the photometric consistency constraint on the FlowNet, optical flow in our CPG stream transitions to “quasi-rigid flow” under the guidance of static scene hypothesis in the PCG stream at the early stage of training. This enables our CGDC and DPC losses in the CPG stream to be effectively applied across the entire image, irrespective of static and dynamic regions. Moreover, as training progresses, photometric loss tends to mislead DepthNet’s predictions. At this time, (16) counteracts the misdirection caused by the photometric loss, ensuring that the DepthNet performs accurately in dynamic regions. We conduct an ablation study in Sect. IV-E to demonstrate the effectiveness of this strategy.
Figure 3: An illustration of the interaction between PCG and CPG streams through the proposed BSCA strategy to address the challenges posed by dynamic objects.
@@
TABLE I: Quantitative comparison with SoTA networks on the KITTI [43], DDAD [29], nuScenes [44] and Waymo Open [45] benchmarks. The best results are shown in bold symbol. The symbols $\uparrow$ and $\downarrow$ indicate that higher and lower values correspond to better performance, respectively. All models are trained with monocular video sequences. The baseline models used for our experiments on each dataset are underlined, respectively.
Dataset Method Year Resolution (pixels) Abs Rel $\downarrow$ Sq Rel $\downarrow$ RMSE $\downarrow$ RMSE log $\downarrow$ $\delta<1.25$uparrow $\delta<1.25^{2}$uparrow $\delta<1.25^{3}$uparrow
KITTI Monodepth2[9] 2019 192 $\times$ 640 0.115 0.903 4.863 0.193 0.877 0.959 0.981
DIFFNet[7] 2021 192 $\times$ 640 0.102 0.764 4.483 0.180 0.890 0.964 0.983
DaCCN[31] 2023 192 $\times$ 640 0.099 0.661 4.316 0.173 0.897 0.967 0.985
Dynamo-Depth[42] 2023 192 $\times$ 640 0.112 0.758 4.505 0.183 0.873 0.959 0.984
Lite-Mono-8M[12] 2023 192 $\times$ 640 0.101 0.729 4.454 0.178 0.897 0.965 0.983
SENSE[46] 2023 192 $\times$ 640 0.104 0.693 4.294 0.177 0.894 0.965 0.984
MonoDiffusion[47] 2024 192 $\times$ 640 0.099 0.702 4.385 0.176 0.899 0.966 0.984
AQUANet[41] 2024 192 $\times$ 640 0.105 0.621 4.227 0.179 0.889 0.964 0.984
DCPI-Depth (Ours) - 192 $\times$ 640 0.095 0.662 4.274 0.170 0.902 0.967 0.985
DDAD Monodepth2[9] 2019 384 $\times$ 640 0.239 12.547 18.392 0.316 0.752 0.899 0.949
DIFFNet[7] 2021 384 $\times$ 640 0.205 12.126 18.461 0.289 0.795 0.916 0.957
SC-Depth[37] 2021 384 $\times$ 640 0.169 3.877 16.290 0.280 0.773 0.905 0.951
Lite-Mono[12] 2023 384 $\times$ 640 0.161 4.451 16.261 0.271 0.802 0.921 0.962
Lite-Mono-8M[12] 2023 384 $\times$ 640 0.175 6.425 16.687 0.272 0.799 0.920 0.961
SC-DepthV3[11] 2023 384 $\times$ 640 0.142 3.031 15.868 0.248 0.813 0.922 0.963
DCPI-Depth (Ours) - 384 $\times$ 640 0.140 2.866 15.786 0.238 0.815 0.929 0.970
nuScenes Monodepth2[9] 2019 288 $\times$ 512 0.425 16.592 10.040 0.402 0.723 0.837 0.887
DIFFNet[7] 2021 288 $\times$ 512 0.228 5.925 8.897 0.290 0.772 0.905 0.950
MonoViT-tiny[8] 2022 288 $\times$ 512 0.412 16.061 10.504 0.385 0.717 0.842 0.898
Lite-Mono[12] 2023 288 $\times$ 512 0.419 15.578 9.807 0.449 0.720 0.831 0.879
Lite-Mono-8M[12] 2023 288 $\times$ 512 0.429 17.058 10.559 0.400 0.709 0.830 0.883
Dynamo-Depth[42] 2023 288 $\times$ 512 0.179 2.118 7.050 0.271 0.787 0.896 0.940
DCPI-Depth (Ours) - 288 $\times$ 512 0.160 1.736 7.194 0.248 0.793 0.921 0.966
Waymo Monodepth2[9] 2019 320 $\times$ 480 0.173 2.731 7.708 0.227 0.797 0.930 0.968
DIFFNet[7] 2021 320 $\times$ 480 0.149 2.082 7.474 0.200 0.838 0.956 0.981
Li et al.[48] 2021 320 $\times$ 480 0.157 1.531 7.090 0.205
Lite-Mono[12] 2023 320 $\times$ 480 0.158 2.305 7.394 0.210 0.816 0.944 0.976
Lite-Mono-8M[12] 2023 320 $\times$ 480 0.154 2.297 7.495 0.209 0.825 0.947 0.975
Dynamo-Depth[42] 2023 320 $\times$ 480 0.116 1.156 6.000 0.166 0.878 0.969 0.989
DCPI-Depth (Ours) - 320 $\times$ 480 0.116 0.963 5.642 0.162 0.872 0.972 0.991
IV Experiments
IV-A Datasets
We conduct our experiments on six public datasets: KITTI[43], DDAD[29], nuScenes [44], Waymo Open [45], Make3D[49], and DIML [50].
For the KITTI[43] dataset, we adopt the Eigen split [5], which comprises 39,180 monocular triplets for training, 4,424 images for validation, and 697 images for testing. For the DDAD[29] dataset, we follow the prior work [11] to split this dataset into a training set of 12,650 images and a test set of 3,950 images in our experiments. For the nuScenes [44] dataset, following [42], we use 79,760 image triplets collected by the front camera for model training, and evaluate the model’s performance on 6,019 front camera images. For the Waymo Open [45] dataset, as in [42], we utilize 76,852 front camera image triplets for training, and 2,216 front camera images for evaluation. For the Make3D[49] and DIML [50] datasets, since neither stereo image pairs nor monocular sequences are provided for unsupervised training, we only use this dataset to quantify the generalizability of models per-trained on the KITTI dataset.
IV-B Experimental Setup
Our experiments are conducted on an NVIDIA RTX 4090 GPU with a batch size of 12. Following [9], we adopt a training approach wherein a snippet of three consecutive video frames is utilized as a training sample. To augment the dataset, random color jitter and horizontal flips are applied to the images during model training. To minimize the loss functions, we employ the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$ and a weight decay of $1\times 10^{-2}$. The learning rate is adjusted using a cosine annealing scheduler with periodic restarts, decaying from $1\times 10^{-4}$ to $5\times 10^{-6}$ over 31 epochs. A decay factor of $\gamma=0.9$ is applied after each cycle to ensure gradual reduction in learning rates. Following [9, 12], the network’s encoder is initialized using pre-trained weights from the ImageNet database [51]. RAFT [15] is employed as the FlowNet in our framework to provide dense correspondence prior. It is pre-trained on the KITTI Flow 2015 [52] dataset, which contains only 200 sets of two consecutive frames and has a small overlap with the KITTI Eigen split [5].
IV-C Evaluation Metrics
Figure 5: Qualitative comparisons between Lite-Mono and our proposed DCPI-Depth on the DDAD [29], nuScenes [44] and Waymo Open [45] datasets.
TABLE II: Quantitative comparison with SoTA networks on the KITTI [43] dataset using the improved KITTI ground truth provided in [53] for testing or higher-resolution images for training. The best results are shown in bold symbol. The symbols $\uparrow$ and $\downarrow$ indicate that higher and lower values correspond to better performance, respectively. “M” denotes training with monocular video sequences, and “MS” denotes training with both monocular video sequences and stereo image pairs. indicates the results achieved using the same weights in Table I for consistency.